OpenStudy (anonymous):

How many real solutions does this system of equations have? x^2 + y^2=25 and x-3y=15

OpenStudy (goformit100):

sIMPLY Solve Equations x^2 + y^2=25 and x-3y=15

OpenStudy (anonymous):

Can you give me a hint of how to start to solve, cause I have no idea?

OpenStudy (abb0t):

You could set it up in the form: \(y=\sqrt{x^2-z^2}\). You will see that you get a easily factorable quadratic that for which you'll get two roots.

OpenStudy (anonymous):

still dont get it

OpenStudy (mertsj):

Do you recognize that the first equation is a circle with center at the origin and radius 5?

OpenStudy (mertsj):



OpenStudy (anonymous):

yes i see now thats a circle

OpenStudy (mertsj):

Now solve the linear equation for y and graph it.

OpenStudy (mertsj):

Can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when i solve for y i get -x/3-5 is that correct

OpenStudy (mertsj):

I got 1/3x - 5

OpenStudy (anonymous):

yea i did it again and thats what i got too

OpenStudy (mertsj):

Can you graph that?

OpenStudy (mertsj):